Compute the total and annual returns on the described investment.
Five years after buying 150 shares of XYZ stock for $60 per share, you sell the stock for $14,400.

Answer by drk(1908)

150 x 60 = 9000 buy in
total returns = (14400-9000) = 5,400
annual return = 5400/5 = 1,080
---
total return as percentage of investment: 5,400/9,000 = 60%
annual return as percentage is: 12%